OpenStudy (anonymous):

a 10 kg object resting on a horizontal frictionless surface is subjected to two forces:F1=30N directed east and F2= 50N directed west. Find (a) its acceleration and (b) displacement after 10 seconds starting from rest.

OpenStudy (anonymous):

a) F2-F1 = total force = 20 N F= ma 20=10a a=.5 m/s^2

OpenStudy (anonymous):

b) v =vo + at v= 0 + .5*10 v= 5 v^2=vo^2+2ad 25 = 0 +2*.5*d d = 25 m

OpenStudy (anonymous):

how did you get .5 for a?

OpenStudy (anonymous):

the 50 N to the right cancels out the 30 to the left leaving you with a total force of 20 to the right then newtons second law of F=ma gives you the acceleration

OpenStudy (anonymous):

20/10=2

OpenStudy (anonymous):

is it okey to use the formula S=Vot+1/2at^2 instead of v^2=vo^2+2ad

OpenStudy (anonymous):

oh yeah i messed that up

OpenStudy (anonymous):

but yes that equation will work in one step

OpenStudy (anonymous):

both give you the right answer

OpenStudy (anonymous):

so you mean tha accelaration would be 2 while s=100 m,is it right?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no problem!